OpenStudy (anonymous):

1.what is an equation in point-slope form for the line perpendicular to y=3x+9 that contains(-6,5)

OpenStudy (whpalmer4):

1. Find the slope of the line \(y = 3x+9\) by recognizing that it is in slope-intercept form, \(y = mx+b\) where \(m\) is the slope 2. Find the slope of a line perpendicular to that: the new slope will be \(\large m_p = -\frac{1}{m}\) 3. Use the point-slope formula \(y-y_1 = m(x-x_1)\) to write the equation of a line that goes through point \((x_1,y_1)\) and has slope \(m\) (note that you will use \(m_p\) here!)